Is it possible to solve a question like this with a TI-84 plus CE

Suppose you are hired by a marketing firm, wanting to know how long a person will watch one of their youtube adds before clicking past it. You are hired to find out how large of a survey/how large of a sample size they will need to answer that question. They want to know with 95% confidence, and they want to know within 1 second. You take the time and collect a small survey of 100 people. You find that the mean is 15 seconds, with a sample standard deviation of 4 seconds. What sample size is necessary to give a 95% confidence interval at the desired level of precision?

  • A. 62
  • B. 2
  • C. 15.68
  • D. 61

If so how?
Answer #1

Solution:

Margin of error = E = 1

= 4

c = 95% = 0.95

= 1 - c = 1 - 0.95 = 0.05

/2 = 0.025

1 - (/2) = 1 - 0.025 = 0.975

Formula for sample size is

Use TI-84 plus CE

Press 2ND and then press VARS to display DISTR menu.

Go to invNorm

invNorm(0.975 , 0 , 1)

= 1.95996398454

1.95996398454 * 4 = 7.83985593816

7.83985593816/1 = 7.83985593816

(7.83985593816)2 = 61.4633411311

Sample size is always rounded to the next whole number

So , answer is 62